OpenStudy (anonymous):

integrate 1/t(t^2-1)^1/2

OpenStudy (anonymous):

A long way would be to allow \(u=t^2-1\). This is awfully ugly and required integration by parts. Though, I'm sure there is a nicer way.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

cn u solve it completely

OpenStudy (anonymous):

I mean, it can be solved completely, but I'm not going to write it all up. Here's the gist: \[ \int_\gamma \frac{dt}{t\sqrt{t^2-1}}=\int_\gamma \frac{du}{(u-1)\sqrt{u}} \]Do integration by parts on the last integral and you should get the answer.

OpenStudy (anonymous):

That should be a \(u+1\), not \(u-1\).

OpenStudy (anonymous):

Made a mistake, the integral is this: \[ \int_\gamma \frac{dt}{t\sqrt{t^2-1}}=\frac{1}{2}\int_\gamma \frac{du}{(u+1)\sqrt{u}} \]

OpenStudy (anonymous):

ok i got it thanks

OpenStudy (anonymous):

Easier method: Let \(u=\sqrt{t^2-1}\)!